# De-mark: Watermark Removal in Large Language Models

### v1

@@ Abstract
Watermarking techniques offer a promising way to identify machine-generated content via embedding covert information into the contents generated from language models (LMs). However, the robustness of the watermarking schemes has not been well explored. In this paper, we present De-mark, an advanced framework designed to remove n-gram-based watermarks effectively. Our method utilizes a novel querying strategy, termed random selection probing, which aids in assessing the strength of the watermark and identifying the red-green list within the n-gram watermark. Experiments on popular LMs, such as Llama3 and ChatGPT, demonstrate the efficiency and effectiveness of De-mark in watermark removal and exploitation tasks.
De-mark: Watermark Removal in Large Language Models
Ruibo Chenthanks: Equal contributions, Yihan Wu, Junfeng Guo, Heng Huang University of Maryland, College Park {rbchen,ywu42}@umd.edu
@@ 1 Introduction
As language models rapidly evolve and their usage becomes more widespread Wang et al. (2024a); Thirunavukarasu et al. (2023); Rubenstein et al. (2023); Wang et al. (2024b); Kasneci et al. (2023), concerns among researchers and regulators about their potential for misuse, particularly in generating harmful content, have intensified. The challenge of distinguishing between human-written and AI-generated text has emerged as a key focus of research. Statistical watermarking (Aaronson, 2022; Kirchenbauer et al., 2023a) presents a promising approach to identifying text produced by sequential LLMs. This method embeds a statistical pattern into the generated text via a pseudo-random generator, which can later be detected through a statistical hypothesis test.
However, recent research Jovanović et al. (2024) highlights that the robustness of existing n-gram watermarking schemes remains insufficiently understood and, in many cases, significantly overestimated. One critical vulnerability lies in watermark stealing & removal, a method that seeks to reverse-engineer the watermarking process. With this method, adversaries can effectively strip or alter the embedded watermark, undermining its intended protection. This poses a serious challenge to the security and integrity of watermarked content, as it exposes a weakness in the watermarking scheme’s ability to withstand tampering.
In this paper, we delve deeper into watermark removal techniques and introduce De-mark, a comprehensive framework for stealing and removing n-gram-based watermarks. Our approach utilizes a querying strategy on the language model to reconstruct the red and green lists and estimate the watermark strength. Unlike prior work Jovanović et al. (2024), which has certain limitations – namely, a) requiring knowledge of the underlying hash function of the n-gram and b) requiring paraphrase tool for watermark removal, which cannot preserve the original LM distribution – our method is more general and provides a theoretical guarantee between the post-removal LM distribution and the original ones. Additionally, De-mark can be adapted for watermark exploitation by imitating the watermarking rule and employing another language model to generate watermarked content.
@@
  • We propose De-mark, a comprehensive watermark framework designed for the removal of n-gram watermarks Kirchenbauer et al. (2023a), which is considered a pivotal approach in the field. Our framework introduces a provable unbiased estimator to assess the strength of watermarks and offers theoretical guarantees about the gap between the original and the post-removal distributions of the language model. Essentially, De-mark operates without requiring prior knowledge of the n-gram watermark parameters.
@@ 2 Related works
Aaronson (2022) introduced the first statistical watermarking algorithm using Gumbel sampling. Building on this, Kirchenbauer et al. (2023a) expanded the framework by categorizing tokens into red and green lists and encouraging the selection of green tokens. Huo et al. (2024) further enhanced the red-green list watermark’s detectability by using learnable networks to optimize the separation of red and green tokens and boosting the green list logits. Zhao et al. (2023) developed the unigram watermark, improving the robustness of statistical watermarking by utilizing one-gram hashing for generating watermark keys. Liu et al. (2023) contributed to watermark robustness by incorporating content semantics as watermark keys. Other researchers, including Christ et al. (2023), Kuditipudi et al. (2023), Hu et al. (2023a), Wu et al. (2023b), Wu et al. (2024), and Chen et al. (2024b) have focused on distortion-free watermarking, aiming to retain the original distribution of the language model.
@@ 3 Preliminary
n-gram watermarking Kirchenbauer et al. (2023a, b) refers the watermarking approaches that use a fixed secret key $\textup{{sk}}_{0}$ and the prefix n-gram $s$ (e.g., $s=\bm{x}_{l-n:l-1}$ for generating $x_{l}$) to form the watermark keys, i.e., $K=\{(\textup{{sk}}_{0},s)\mid s\in\mathcal{V}^{n}\}$, where $\mathcal{V}^{n}$ represents the set of all n-grams with token set $V$. After that, the watermark keys are used to divide the token set into a red list and a green list, with the logits of the green tokens being increased by $\delta$ (a.k.a. watermark strength). Specifically, given an initial token probability $P(t)$, the watermarked probability for the token, denoted by $P_{W}(t)$, is:
$P_{W}(t)=\begin{cases}&\frac{P(t)}{\sum_{t^{\prime}\in\textrm{red}}P(t^{\prime})+\sum_{t^{\prime}\in\textrm{green}}e^{\delta}P(t^{\prime})},t\in\textrm{red};\\ &\frac{e^{\delta}P(t)}{\sum_{t^{\prime}\in\textrm{red}}P(t^{\prime})+\sum_{t^{\prime}\in\textrm{green}}e^{\delta}P(t^{\prime})},t\in\textrm{green},\end{cases}$ (1)
@@ 3.1 Threat models & settings
We select No access scenario (D0) Jovanović et al. (2024), where the watermark detector remains completely private. This setup, which has been widely considered in prior research, imposes stricter limitations on the attacker.
@@
We select Unavailable base responses setting (B0) Jovanović et al. (2024), where the attacker does not have access to the non-watermarked models.
@@ 4.1.1 Calculate relative probability ratio
Define $V:=\{t_{1},\dots,t_{N}\}$ as the vocabulary or set of tokens of a language model, $T=\{T_{1},T_{2},\dots,T_{m}\}\subset V$, and $P(T_{i}|\bm{x},T_{i},T_{j})$ the LM probability of token $T_{i}$ generated with random selection probing. We first prompt an LM to randomly output two tokens $T_{i}$ and $T_{j}$ with equal probability. In the ideal case $P(T_{i}|\bm{x},T_{i},T_{j})=P(T_{j}|\bm{x},T_{i},T_{j})=0.5$. However, in practice $P(T_{i}|\bm{x},T_{i},T_{j})$ and $P(T_{j}|\bm{x},T_{i},T_{j})$ can not be perfectly $0.5$. Thus we assume a noisy case: $P(T_{i}|\bm{x},T_{i},T_{j})=0.5+\epsilon$, $P(T_{j}|\bm{x},T_{i},T_{j})=0.5-\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. We define relative probability ratio, denoted as $R_{i,j}:=\log\left(\frac{P(T_{i}|\bm{x},T_{i},T_{j})}{P(T_{j}|\bm{x},T_{i},T_{j})}\right)$. The intuition behind $R_{i,j}$ is as follows: in n-gram watermarking, if one token belongs to the red list and the other to the green list, and their original probabilities are equal, then following Equation 1, $\log\left(\frac{P_{W}(T_{i}|\bm{x},T_{i},T_{j})}{P_{W}(T_{j}|\bm{x},T_{i},T_{j})}\right)$ will be either $\delta$ or $-\delta$. Conversely, if both tokens are from the red or green list, $\log\left(\frac{P_{W}(T_{i}|\bm{x},T_{i},T_{j})}{P_{W}(T_{j}|\bm{x},T_{i},T_{j})}\right)$ will be 0. This allows us to efficiently estimate the value of $\delta$ using $R_{i,j}$. In the following theorem, we show $R_{i,j}$ is an unbiased estimator of $\delta$ (or $-\delta$) if $T_{i}$ and $T_{j}$ are not in the same red or green list.
@@
Note that LLMs are sensitive to the relative positioning of tokens (Lee et al., 2023; Wang et al., 2023a; Pezeshkpour and Hruschka, 2023). In calculating $P(T_{i}|\bm{x},T_{i},T_{j})$ and $P(T_{j}|\bm{x},T_{i},T_{j})$, we adjust the positions of $T_{i}$ and $T_{j}$ during generation. The average probability from these swapped positions is then used as the final probability for each token. Detailed formulations for $F$ and the specific prompts employed are provided in Appendix C.1.
@@ 4.1.2 Calculate token score
4:     Generate token score $\bm{s}^{h^{\prime}}$ using $\bm{x}[-h^{\prime}:]$, $T$ with Alg.5
@@ 4.1.3 Identify the prefix n-gram length $h$
5:     Generate token score $\bm{s}^{k}$ using $\bm{x}$, $T$ with Alg.5, generate relative probability ratio $R^{k}$ using Alg.4
@@ 4.1.4 Identify watermark strength $\delta$
2:Generate relative probability ratio $R$ with Alg.4
@@
TPR@FPR= median p-value GPT score TPR@FPR= median p-value GPT score
0.1% 0.01% 0.001% 0.1% 0.01% 0.001%
@@
TPR@FPR= median p-value GPT score TPR@FPR= median p-value GPT score
0.1% 0.01% 0.001% 0.1% 0.01% 0.001%
@@ 4.2 De-mark applications
After identifying the watermark strength $\hat{\delta}$ and the red-green lists $\hat{R}$ and $\hat{G}$, we can proceed with removing the watermark from the model. Note that our watermark removal approach differs from the method described in Jovanović et al. (2024). While they employ paraphrasing techniques, which alter the original LM distribution, our strategy is designed to preserve the original LM distribution as much as possible. Given the watermarked distribution $P_{W}$, the following rule is applied for watermark removal. Denote the distribution after removal by $P_{R}$:
$P_{R}(t)=\begin{cases}&\frac{P_{W}(t)}{\sum_{t\in\hat{R}}P_{W}(t)+\sum_{t\in\hat{G}}e^{-\eta\hat{\delta}}P_{W}(t)},\quad t\in\hat{R};\\ &\frac{e^{-\eta\hat{\delta}}P_{W}(t)}{\sum_{t\in\hat{R}}P_{W}(t)+\sum_{t\in\hat{G}}e^{-\eta\hat{\delta}}P_{W}(t)},\quad t\in\hat{G},\end{cases}$ (2)
@@ 5 Experiments
Our experimental evaluation is structured into three main parts. In the first part, we present the primary results of watermark removal and its exploitation. In the second part, we validate the effectiveness of De-mark in identifying key parameters, including the prefix n-gram length $h$, the watermark strength $\delta$, and the red-green list, under both gray-box setting, where the top-k probability is known (L1), and black-box setting, where the top-k probability is unavailable (L0). Finally, in the case study, we demonstrate the practical application of De-mark on ChatGPT. The detailed experimental settings are in Appendix C.2. Additional experimental results and ablation study can be found in Appendix D and E.
@@ 5.1 Watermark removal & exploitation
We evaluate De-mark on both watermark removal and watermark exploitation tasks. For watermark removal, we apply n-gram watermarking Kirchenbauer et al. (2023a) to the generated content, using an unknown prefix n-gram length $h$, watermark strength $\delta$, and a specified red-green token list. We report the true positive rate at theoretical false positive rates (TPR@FPR) of 0.1%, 0.01%, 0.001% FPR and the median p-value. For watermark exploitation, we use a language model along with estimated watermarking rules to rephrase the original content and embed watermarks.

### v2

@@ Abstract
Watermarking techniques offer a promising way to identify machine-generated content via embedding covert information into the contents generated from language models (LMs). However, the robustness of the watermarking schemes has not been well explored. In this paper, we present De-mark, an advanced framework designed to remove n-gram-based watermarks effectively. Our method utilizes a novel querying strategy, termed random selection probing, which aids in assessing the strength of the watermark and identifying the red-green list within the n-gram watermark. Experiments on popular LMs, such as Llama3 and ChatGPT, demonstrate the efficiency and effectiveness of De-mark in watermark removal and exploitation tasks. Our code is available at https://github.com/RayRuiboChen/De-mark.
Machine Learning, ICML
@@ 1 Introduction
As language models rapidly evolve and their usage becomes more widespread (Wang et al., 2024a; Thirunavukarasu et al., 2023; Rubenstein et al., 2023; Wang et al., 2024b; Kasneci et al., 2023; Wang et al., 2023a; Chen et al., 2024b), concerns among researchers and regulators about their potential for misuse, particularly in generating harmful content, have intensified. The challenge of distinguishing between human-written and AI-generated text has emerged as a key focus of research. Statistical watermarking (Aaronson, 2022; Kirchenbauer et al., 2023a) presents a promising approach to identifying text produced by sequential LLMs. This method embeds a statistical pattern into the generated text via a pseudo-random generator, which can later be detected through a statistical hypothesis test.
However, recent research (Jovanović et al., 2024) highlights that the robustness of existing n-gram watermarking schemes remains insufficiently understood and, in many cases, significantly overestimated. One critical vulnerability lies in watermark stealing & removal, a method that seeks to reverse-engineer the watermarking process. With this method, adversaries can effectively strip or alter the embedded watermark, undermining its intended protection. This poses a serious challenge to the security and integrity of watermarked content, as it exposes a weakness in the watermarking scheme’s ability to withstand tampering.
In this paper, we delve deeper into watermark removal techniques and introduce De-mark, a comprehensive framework for stealing and removing n-gram-based watermarks. Our approach utilizes a querying strategy on the language model to reconstruct the red and green lists and estimate the watermark strength. Unlike prior work (Jovanović et al., 2024), which has certain limitations – namely, a) requiring knowledge of the underlying hash function of the n-gram and b) requiring paraphrase tool for watermark removal, which cannot preserve the original LM distribution – our method is more general and provides a theoretical guarantee between the post-removal LM distribution and the original ones. Additionally, De-mark can be adapted for watermark exploitation by imitating the watermarking rule and employing another language model to generate watermarked content.
@@
  • We propose De-mark, a comprehensive watermark framework designed for the removal of n-gram watermarks (Kirchenbauer et al., 2023a), which is considered a pivotal approach in the field. Our framework introduces a provable unbiased estimator to assess the strength of watermarks and offers theoretical guarantees about the gap between the original and the post-removal distributions of the language model. Essentially, De-mark operates without requiring prior knowledge of the n-gram watermark parameters.
@@ 2 Related works
Aaronson (2022) introduced the first statistical watermarking algorithm using Gumbel sampling. Building on this, Kirchenbauer et al. (2023a) expanded the framework by categorizing tokens into red and green lists and encouraging the selection of green tokens. Huo et al. (2024) further enhanced the red-green list watermark’s detectability by using learnable networks to optimize the separation of red and green tokens and boosting the green list logits. Zhao et al. (2023) developed the unigram watermark, improving the robustness of statistical watermarking by utilizing one-gram hashing for generating watermark keys. Liu et al. (2023) contributed to watermark robustness by incorporating content semantics as watermark keys. Other researchers, including Christ et al. (2023), Kuditipudi et al. (2023), Hu et al. (2023), Wu et al. (2023b), Wu et al. (2024), and Chen et al. (2024c) have focused on distortion-free watermarking, aiming to retain the original distribution of the language model.
@@
Additionally, Gloaguen et al. (2024) and Liu et al. (2024a) investigate methods for detecting language model watermarks, whereas Liu et al. (2024b) demonstrates that image watermarks can be removed.
We also acknowledge the concurrent research by Pang et al. (2024) and Zhang et al. (2024), which also explores watermark stealing. However, The experimental settings of them differ significantly from ours. For example, Pang et al. (2024) focused on detector-available scenarios, while we address the more challenging detector-unavailable setting. Zhang et al. (2024) focused only on Unigram watermark and is very difficult to adapt to the n-gram setting we are working on. Thus, we are unable to provide direct comparisons.
@@ 3 Preliminary
n-gram watermarking (Kirchenbauer et al., 2023a, b) refers the watermarking approaches that use a fixed secret key $\textup{{sk}}_{0}$ and the prefix n-gram $s$ (e.g., $s=\bm{x}_{l-n:l-1}$ for generating $x_{l}$) to form the watermark keys, i.e., $K=\{(\textup{{sk}}_{0},s)\mid s\in\mathcal{V}^{n}\}$, where $\mathcal{V}^{n}$ represents the set of all n-grams with token set $V$. After that, the watermark keys are used to divide the token set into a red list and a green list, with the logits of the green tokens being increased by $\delta$ (a.k.a. watermark strength). Specifically, given an initial token probability $P(t)$, the watermarked probability for the token, denoted by $P_{W}(t)$, is:
$\small P_{W}(t)=\begin{cases}&\frac{P(t)}{\sum_{t^{\prime}\in\textrm{red}}P(t^{\prime})+\sum_{t^{\prime}\in\textrm{green}}e^{\delta}P(t^{\prime})},t\in\textrm{red};\\ &\frac{e^{\delta}P(t)}{\sum_{t^{\prime}\in\textrm{red}}P(t^{\prime})+\sum_{t^{\prime}\in\textrm{green}}e^{\delta}P(t^{\prime})},t\in\textrm{green},\end{cases}$ (1)
@@ 3.1 Threat models & settings
We select No access scenario (D0) (Jovanović et al., 2024), where the watermark detector remains completely private. This setup, which has been widely considered in prior research, imposes stricter limitations on the attacker.
@@
We select Unavailable base responses setting (B0) (Jovanović et al., 2024), where the attacker does not have access to the non-watermarked models.
@@ 4.1 Main algorithms
Algorithm 1 Calculate relative probability ratio
2:Output: relative ratio matrix $R\in\mathbb{R}^{m\times m}$
3:Initialize $R=0$
4:for $i=1,\dots,m$ do
5:     for $j=i+1,\dots,m$ do
6:         $P(T_{i}|\bm{x},T_{i},T_{j}),P(T_{j}|\bm{x},T_{i},T_{j})=F(\bm{x},T_{i},T_{j})$
7:         $R_{i,j}=log(\frac{P(T_{i}|\bm{x},T_{i},T_{j})}{P(T_{j}|\bm{x},T_{i},T_{j})})$
8:         $R_{j,i}=log(\frac{P(T_{j}|\bm{x},T_{i},T_{j})}{P(T_{i}|\bm{x},T_{i},T_{j})})$
9:Return: $R$
@@ 4.1.1 Calculate relative probability ratio
Define $V:=\{t_{1},\dots,t_{N}\}$ as the vocabulary or set of tokens of a language model, $T=\{T_{1},T_{2},\dots,T_{m}\}\subset V$, and $P(T_{i}|\bm{x},T_{i},T_{j})$ the LM probability of token $T_{i}$ generated with random selection probing. We first prompt an LM to randomly output two tokens $T_{i}$ and $T_{j}$ with equal probability. In the ideal case $P(T_{i}|\bm{x},T_{i},T_{j})=P(T_{j}|\bm{x},T_{i},T_{j})=0.5$. However, in practice $P(T_{i}|\bm{x},T_{i},T_{j})$ and $P(T_{j}|\bm{x},T_{i},T_{j})$ can not be perfectly $0.5$. Thus we assume a noisy case: $P(T_{i}|\bm{x},T_{i},T_{j})=0.5+\epsilon$, $P(T_{j}|\bm{x},T_{i},T_{j})=0.5-\epsilon$, where $\epsilon$ is symmetrically distributed, i.e., $p(\epsilon=\epsilon_{0})=p(\epsilon=-\epsilon_{0}),\forall\epsilon_{0}\in\mathbb{R}$ We define relative probability ratio, denoted as $R_{i,j}:=\log\left(\frac{P(T_{i}|\bm{x},T_{i},T_{j})}{P(T_{j}|\bm{x},T_{i},T_{j})}\right)$. The intuition behind $R_{i,j}$ is as follows: in n-gram watermarking, if one token belongs to the red list and the other to the green list, and their original probabilities are equal, then following Equation 1, $\log\left(\frac{P_{W}(T_{i}|\bm{x},T_{i},T_{j})}{P_{W}(T_{j}|\bm{x},T_{i},T_{j})}\right)$ will be either $\delta$ or $-\delta$. Conversely, if both tokens are from the red or green list, $\log\left(\frac{P_{W}(T_{i}|\bm{x},T_{i},T_{j})}{P_{W}(T_{j}|\bm{x},T_{i},T_{j})}\right)$ will be 0. This allows us to efficiently estimate the value of $\delta$ using $R_{i,j}$. In the following theorem, we show $R_{i,j}$ is an unbiased estimator of $\delta$ (or $-\delta$) if $T_{i}$ and $T_{j}$ are not in the same red or green list.
@@
Note that LLMs are sensitive to the relative positioning of tokens (Lee et al., 2023; Wang et al., 2023b; Pezeshkpour & Hruschka, 2023). In calculating $P(T_{i}|\bm{x},T_{i},T_{j})$ and $P(T_{j}|\bm{x},T_{i},T_{j})$, we adjust the positions of $T_{i}$ and $T_{j}$ during generation. The average probability from these swapped positions is then used as the final probability for each token. Detailed formulations for $F$ and the specific prompts employed are provided in Appendix B.1.
Algorithm 2 Calculate token score
2:Output: token scores $\bm{s}\in\mathbb{R}^{m}$
3:Initialize $\bm{s}=0$
4:Generate $R$ using $\bm{x}$, $T$ with Alg.1
5:for $i=1,\dots,m$ do
6:     for $j=1,\dots,m$ do
7:         if $\alpha_{1}<R_{i,j}<\alpha_{2}$ then
8:              $s_{i}+=1$
9:         else if $-\alpha_{2}<R_{i,j}<-\alpha_{1}$ then
10:              $s_{i}-=1$
11:Return: $\bm{s}$
@@ 4.1.2 Calculate token score
4:     Generate token score $\bm{s}^{h^{\prime}}$ using $\bm{x}[-h^{\prime}:]$, $T$ with Alg.2
@@ 4.1.3 Identify the prefix n-gram length $h$
5:     Generate token score $\bm{s}^{k}$ using $\bm{x}$, $T$ with Alg.2, generate relative probability ratio $R^{k}$ using Alg.1
@@ 4.1.4 Identify watermark strength $\delta$
2:Generate relative probability ratio $R$ with Alg.1
@@
TPR@FPR= median $\uparrow$ GPT $\uparrow$ TPR@FPR= median $\uparrow$ GPT $\uparrow$
0.1$\downarrow$ 0.01$\downarrow$ 0.001$\downarrow$ 0.1$\downarrow$ 0.01$\downarrow$ 0.001$\downarrow$
@@
TPR@FPR= median $\downarrow$ GPT $\uparrow$ TPR@FPR= median $\downarrow$ GPT $\uparrow$
0.1$\uparrow$ 0.01$\uparrow$ 0.001$\uparrow$ 0.1$\uparrow$ 0.01$\uparrow$ 0.001$\uparrow$
@@ 4.2 De-mark applications
After identifying the watermark strength $\hat{\delta}$ and the red-green lists $\hat{R}$ and $\hat{G}$, we can proceed with removing the watermark from the model. Note that our watermark removal approach differs from the method described in (Jovanović et al., 2024). While they employ paraphrasing techniques, which alter the original LM distribution, our strategy is designed to preserve the original LM distribution as much as possible. Given the watermarked distribution $P_{W}$, the following rule is applied for watermark removal. Denote the distribution after removal by $P_{R}$:
$\small P_{R}(t)=\begin{cases}&\frac{P_{W}(t)}{\sum_{t\in\hat{R}}P_{W}(t)+\sum_{t\in\hat{G}}e^{-\eta\hat{\delta}}P_{W}(t)},\quad t\in\hat{R};\\ &\frac{e^{-\eta\hat{\delta}}P_{W}(t)}{\sum_{t\in\hat{R}}P_{W}(t)+\sum_{t\in\hat{G}}e^{-\eta\hat{\delta}}P_{W}(t)},\quad t\in\hat{G},\end{cases}$ (2)
@@ 5 Experiments
Our experimental evaluation is structured into three main parts. In the first part, we present the primary results of watermark removal and its exploitation. In the second part, we validate the effectiveness of De-mark in identifying key parameters, including the prefix n-gram length $h$, the watermark strength $\delta$, and the red-green list, under both gray-box setting, where the top-k probability is known (L1), and black-box setting, where the top-k probability is unavailable (L0). Finally, in the case study, we demonstrate the practical application of De-mark on ChatGPT. The detailed experimental settings are in Appendix B.2. Additional experimental results and ablation study can be found in Appendix C and D.
Settings.
For watermark stealing and watermark removal, we use Llama3.1-8B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Llama3.2-3B for our experiments. For watermark exploitation, we choose Llama3.2-1B as the attacker LLM, and the original watermarked LLM is Llama3.2-3B. In gray-box setting (L1), we use top-20 log-probabilities. We generate 300 tokens for each prompt, and suppress the EOS token following (Kirchenbauer et al., 2023a).
Datasets.
Text quality measure.
@@ 5.1 Watermark removal & exploitation
We evaluate De-mark on both watermark removal and watermark exploitation tasks. For watermark removal, we apply n-gram watermarking (Kirchenbauer et al., 2023a) to the generated content, using an unknown prefix n-gram length $h$, watermark strength $\delta$, and a specified red-green token list. We report the true positive rate at theoretical false positive rates (TPR@FPR) of 0.1%, 0.01%, 0.001% FPR and the median p-value. For watermark exploitation, we use a language model along with estimated watermarking rules to rephrase the original content and embed watermarks.